OpenStudy (erinweeks):

I Tried Doing This On My Own & I Figured Out Its Not A But I Cant Get Any Further.. Solve the system by the addition method.

OpenStudy (erinweeks):

A. {( 1, 2), ( -1, 2), ( 1, -2), ( -1, -2)} B. {( 1, 2), ( -1, -2)} C. {( 2, 1), ( -2, 1), ( 2, -1), ( -2, -1)} D. {( 2, 1), ( -2, -1)}

OpenStudy (erinweeks):

@ganeshie8 ?

ganeshie8 (ganeshie8):

where is the system

OpenStudy (erinweeks):

Sorry lol. x2 - 3y2 = 1 3x2 + 3y2 = 15

ganeshie8 (ganeshie8):

it says use addition method, so its going to be easy.

ganeshie8 (ganeshie8):

oly additions we will be doing..

OpenStudy (erinweeks):

i figured out its not a .. but i think im making it more harder than it is.

ganeshie8 (ganeshie8):

lets put the equations in row, and ADD them \(vertically\) x2 - 3y2 = 1 3x2 + 3y2 = 15 ----------------- 4x2 + 0 = 16

ganeshie8 (ganeshie8):

4x2 + 0 = 16 x2 = 4 x = \(\pm\) 2

OpenStudy (erinweeks):

thats kinda were i confused myself at lol

ganeshie8 (ganeshie8):

haa it can be confusing, we have two values for x, x = +2, x = -2

OpenStudy (erinweeks):

alright and what about 1s though?

ganeshie8 (ganeshie8):

lets solve y, when x = +2, put x = 2 in first equation, x2 - 3y2 = 1 2^2 - 3y2 = 1 4 - 3y2 = 1 3 = 3y2 y = \(\pm\) 1

ganeshie8 (ganeshie8):

so our solutions are x = \(\pm\)2, y = \(\pm\) 1

ganeshie8 (ganeshie8):

it can be written as :- {(2, 1), (-2, 1), (2, -1), (-2, -1) }

OpenStudy (erinweeks):

so as C?

ganeshie8 (ganeshie8):

Yes